Company Z provides financial services for its customers. They have a debt of $25 million of which they pay $1.8 million per annum in interest expense. They have $105 million in common stock at 6%. The company’s Beta is 1.1 and the risk-free rate is 2.5%. The tax rate is 40%. What is the cost of debt? What is the cost of equity? What is the WACC?
Answer #1

Pre tax cost of debt = Interest/Debt = 1.8/25 = 7.2%

Post tax cost of debt = 7.2%*(1-40%) = 4.32%

Cost of equity = 6% (given)

Wacc = Sum of costs*weights

= 4.32%*25/(25+105) +6%*105/(25+105)

= 5.68%